Find the slope and y-intercept of the line in the equation y=3/4x-2

I would really appreciate any help I can get with this problem, Ijust dont understand this slope stuff at all. Please help me!!!!

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

y=3/4x-2
is already in "slope-intercept" form of:
y=mx+b
where
m is the slope
b is the y-intercept at (0,b)
.
Your equation:
y=3/4x-2
can be written as:
y=(3/4)x+(-2)
so, from "inspection" then we see that the
slope is 3/4
y-intercept is at (0, -2)

Answer by Alan3354(69443)

Find the slope and y-intercept of the line in the equation y=3/4x-2
-----------
Slope is the coefficient of x.
m = 3/4
-----
the y-int = -2